# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ Appendix D Lagrangian Relaxation of Projection Lemma: Proof of Theorem 2
    Even if the set $H(\tilde{\mathbf{A}})\cap C$ has empty interior or some component if $\tilde{\mathbf{A}}$ is identically zero, by the last statement of the Theorem 10 we can say for any closed convex set Q of $H(\tilde{\mathbf{A}})\cap C$ we can design the function $\{\tilde{\mathbf{A}}^{r}\}$ in such a way that $\lim_{r\rightarrow\infty}H(\tilde{\mathbf{A}}^{r})\cap C$ includes $Q$.
As the convergence of $\tilde{\mathbf{A}}_{t}$ is guaranteed now asymptotically, we can guaranty convergence of the following limit $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$.
Statement (b) : Proof of Uniqueness of ${\boldsymbol{\lambda}}^{\star}$ Here, we prove if there exists a confusing instance ${\boldsymbol{\lambda}}^{*}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ which uniquely minimises the the function $\mathcal{D}(.)$ defined as
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\mathrel{\triangleq}\inf_{\boldsymbol{l}\in\mathcal{L}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}$
@@
Thus, we conclude that for a strictly convex function f(x) with supp(x) being convex as well, the set of minimisers is either empty or singleton. Then, we can say ${\boldsymbol{\lambda}}^{\star}$ uniquely minimizes the KL, or say $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$. Let us now once again remind the definition of perturbed alt-set $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})\mathrel{\triangleq}\{{\boldsymbol{\lambda}}\in\mathbb{D}:\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}{\boldsymbol{\lambda}}^{T}(\hat{\boldsymbol{\pi}}^{\star}-\boldsymbol{\pi})<r\}$. Let us denote $\nu(\hat{\boldsymbol{\pi}}^{\star})$ as the neighborhood of $\hat{\boldsymbol{\pi}}^{\star}$. Any $^{\boldsymbol{\pi}}\in\hat{\mathcal{F}}$ is called a neighbor of $\hat{\boldsymbol{\pi}}^{\star}$, if it is an extreme point of $\hat{\mathcal{F}}$ and shares (K-1) active constraints with $\hat{\boldsymbol{\pi}}^{\star}$. Then, we can decompose the perturbed alt-set as $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})=\bigcup_{{\boldsymbol{\pi}}\in\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})}\left\{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-{\boldsymbol{\pi}})<0\right\}$, which is a union of half-spaces for each neighbor. From this decomposition we can observe that $\hat{\boldsymbol{\pi}}^{\star}$ is not the r-optimal policy for ${\boldsymbol{\lambda}}$, i.e, $\{\exists{\boldsymbol{\pi}}^{\prime}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi}):{\boldsymbol{\lambda}}^{T}(\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}-{\boldsymbol{\pi}}^{\prime})<0\}$. Then, it follows similar argument in Carlsson et al., (2024) to argue that the most confusing instance w.r.t ${\boldsymbol{\mu}}$ lies in the boundary of the normal cone, which lands us to Proposition 1.
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=$ $\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-\boldsymbol{\pi}^{{}^{\prime}})=0}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}$
$=$ $\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-\boldsymbol{\pi}^{{}^{\prime}})=0}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}$
Statement (c): Convergence of the sequence $\{\hat{{\boldsymbol{\lambda}}}_{n}\}_{n\in\mathbb{N}}$. In known constraint setting the agent has access to $\mathcal{F}$. That means there is the actual sequence $\{{\boldsymbol{\lambda}}_{n}\}_{n\in\mathbb{N}}$ for which ${\boldsymbol{\lambda}}_{n}\to{\boldsymbol{\lambda}}^{*}$ as $n\to\infty$ since $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ is convex and continuous on $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. But in this setting we try to estimate $\mathcal{F}$ as ^$\hat{\mathcal{F}}_{n}$ at each time step $n\in\mathbb{N}$. So there exists the $\{\hat{{\boldsymbol{\lambda}}}_{n}\}_{n\in\mathbb{N}}$ such that $\hat{{\boldsymbol{\lambda}}}_{n}\in\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ and we have to ensure it converges to the unique optimal ${\boldsymbol{\lambda}}^{*}$ i.e ${\boldsymbol{\lambda}}^{\star}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})\subseteq\lim_{n\rightarrow\infty}\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ implies $\{\hat{{\boldsymbol{\lambda}}}_{n}\}\to{\boldsymbol{\lambda}}^{*}$ as $n\to\infty$
@@
    ^$\hat{\mathcal{F}}_{n}$ for any $n\in\mathbb{N}$ is a superset of $\mathcal{F}$ due to the pessimistic choice of $\mathbf{A}$.
@@
    ^$\hat{\mathcal{F}}_{n}$ is a non-empty compact subset of $\Delta_{K}$ and $\lim_{n\to\infty}\hat{\mathcal{F}}_{n}=\mathcal{F}$.
@@
    $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is a closed convex set and it also is a superset of the real alt-set $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$.
Leveraging these properties we claim that for any ${\boldsymbol{\mu}}\in\mathcal{D}$, $\lim_{n\to\infty}\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})=\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Since we have already proven uniqueness of ${\boldsymbol{\lambda}}$ in statement 2, we say $\hat{{\boldsymbol{\lambda}}}_{n}$ uniquely minimises $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Now from the $(\epsilon,\delta)$-definition of limits we say if $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is an $\epsilon$-cover of $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ for $\epsilon>0$, then $|\hat{{\boldsymbol{\lambda}}}_{n}-{\boldsymbol{\lambda}}^{*}|\leq\delta$ for $\delta>0$ sufficiently small. It implies for a sequence of $\{\hat{\lambda_{n}}\}_{n\in\mathbb{N}}$ we claim $\lim_{n\to\infty}\hat{{\boldsymbol{\lambda}}}_{n}={\boldsymbol{\lambda}}^{*}$ i.e the sequence convergence. Therefore we conclude by the statement itself
@@ Appendix E Characterization of the unique optimal policy: Proof of Theorem 3
Statement 1: Convexity of feasible space and optimal set function. Let us first analyse the properties of $\hat{\mathcal{F}}$. For any two member of $\boldsymbol{\omega}_{1},\boldsymbol{\omega}_{2}\in\hat{\mathcal{F}}$ satisfying $\tilde{\mathbf{A}}\boldsymbol{\omega}_{1}\leq 0$ and $\tilde{\mathbf{A}}\boldsymbol{\omega}_{2}\leq 0$, their convex combination for any $\alpha\in[0,1]$,
$\tilde{\mathbf{A}}\left(\alpha\boldsymbol{\omega}_{1}+(1-\alpha)\boldsymbol{\omega}_{2}\right)=\alpha\tilde{\mathbf{A}}\boldsymbol{\omega}_{1}+(1-\alpha)\tilde{\mathbf{A}}\boldsymbol{\omega}_{2}\leq 0$
@@
(a) The complement of $\hat{\mathcal{F}}$, ^$\hat{\mathcal{F}}^{c}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}:\tilde{\mathbf{A}}\boldsymbol{\pi}>0\}$ is an open set. (b) we have already proven the limit of $\hat{\mathcal{F}}$ to be $\mathcal{F}$ which is always contained by $\hat{\mathcal{F}}$.
@@
Statement 2: Continuity of limit. We have already proven in Section D that $\lim_{t\to\infty}\hat{\mathcal{F}}_{t}\to\mathcal{F}$. Also by convexity of KL and CLT we claim $\hat{{\boldsymbol{\mu}}}_{t}\to{\boldsymbol{\mu}}$ as $t\to\infty$ and since $\boldsymbol{\omega}$ is linear in $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}}_{t})$ it will converge to $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$ as $t\to\infty$, also due to convergence of $\hat{{\boldsymbol{\mu}}}_{t}$. Then we can say that the limiting value is same as the value if we plug in the limits in $\mathcal{D}$ i.e $\lim_{t\to\infty}\mathcal{D}(\boldsymbol{\omega},\hat{{\boldsymbol{\mu}}}_{t},\hat{\mathcal{F}}_{t})=\mathcal{D}(\boldsymbol{\omega}^{*}({\boldsymbol{\mu}}),{\boldsymbol{\mu}},\mathcal{F})$. So we ensure the continuity of $\lim_{t\rightarrow\infty}\mathcal{D}(\boldsymbol{\omega},\hat{{\boldsymbol{\mu}}}_{t},\hat{\mathcal{F}})$.
@@
Statement 4: Upper hemi-continuity of optimal allocation function. We refer to Magureanu et al., (2014) (see Lemma 12) for this proof. We denote $Q(\tilde{\mathbf{A}}^{\prime})\mathrel{\triangleq}\lim_{\hat{\mathcal{F}}\to\mathcal{F}}\max_{\boldsymbol{\omega}\in\Delta_{K}}\bigg{\{}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}^{\prime\prime}\boldsymbol{\omega}\bigg{|}\tilde{\mathbf{A}}^{\prime\prime}\boldsymbol{\omega}\leq 0,\boldsymbol{\omega}_{a}\geq 0\forall i\in[{K}]\bigg{\}}=\boldsymbol{\omega}({\boldsymbol{\mu}})$ where $\tilde{\mathbf{A}}^{\prime\prime}\in\mathbb{R}^{{K}\times{K}}$ is the rank-1 update of $\tilde{\mathbf{A}}^{\prime}$ which is a sub-matrix of $\tilde{\mathbf{A}}$ with ${K}$ number of active constraints. We define limiting set as
$Q^{\star}(\tilde{\mathbf{A}}^{\prime\prime})=\bigg{\{}\boldsymbol{\omega}:\lim_{\hat{\mathcal{F}}\to\mathcal{F}}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)=Q(\tilde{\mathbf{A}}^{\prime\prime})\bigg{|}\tilde{\mathbf{A}}^{\prime\prime}\boldsymbol{\omega}\leq 0,\boldsymbol{\omega}_{a}\geq 0\forall i\in[{K}]\bigg{\}}=\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$
@@ Appendix F Lagrangian Lower Bound for Gaussians: Proof of Theorem 4
$T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})=\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{2\sigma^{2}\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}\,.$
@@
We start the proof by the definition of $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ as per Equation (9)
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ $=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\left\{\sum_{a=1}^{k}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a})-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
@@
The Lagrangian formulation of $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ is written as
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\lambda}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$
@@
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\gamma\in\mathbb{R}_{+}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\sum_{a=1}^{K}\boldsymbol{\omega}_{a}\frac{({\boldsymbol{\mu}}_{a}-{\boldsymbol{\lambda}}_{a})^{2}}{2\sigma^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\lambda}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$ (11)
@@
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\gamma\in\mathbb{R}_{+}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{-\sum_{a=1}^{K}\gamma^{2}\frac{\boldsymbol{v}_{a}^{2}\sigma^{2}}{2\boldsymbol{\omega}_{a}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\mu}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$ (12)
@@
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ $=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{(r-{\boldsymbol{\mu}}^{\top}v)^{2}}{2\sigma^{2}\sum_{a=1}^{K}\frac{\boldsymbol{v}_{a}^{2}}{\boldsymbol{\omega}_{a}}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$
$=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{1}{2\sigma^{2}}\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\right\}$

### v2

@@ Appendix D Lagrangian Relaxation of Projection Lemma: Proof of Theorem 2
    Even if the set $H(\tilde{A})\cap C$ has empty interior or some component if $\tilde{A}$ is identically zero, by the last statement of the Theorem 10 we can say for any closed convex set Q of $H(\tilde{A})\cap C$ we can design the function $\{\tilde{A}^{r}\}$ in such a way that $\lim_{r\rightarrow\infty}H(\tilde{A}^{r})\cap C$ includes $Q$.
As the convergence of $\tilde{A}_{t}$ is guaranteed now asymptotically, we can guaranty convergence of the following limit $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$.
Statement (b) : Proof of Uniqueness of $\boldsymbol{\lambda}$$^{\star}$ Here, we prove if there exists a confusing instance ${\boldsymbol{\lambda}}^{*}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$ which uniquely minimises the the function $\mathcal{D}(.)$ defined as
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\mathrel{\triangleq}\inf_{\boldsymbol{l}\in\mathcal{L}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}$
@@
[…]
@@
[…]
$=$ $\min_{\boldsymbol{l}\in\mathcal{L}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-\boldsymbol{\pi}^{{}^{\prime}})=0}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}$
[…]
@@
    $\hat{\mathcal{F}}_{n}$ for any $n\in\mathbb{N}$ is a superset of $\mathcal{F}$ due to the optimistic choice of $A$.
@@
    $\hat{\mathcal{F}}_{n}$ is a non-empty compact subset of $\Delta_{K}$ and $\lim_{n\to\infty}\hat{\mathcal{F}}_{n}=\mathcal{F}$.
@@
    $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is a closed convex set and it also is a superset of the real alt-set $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$.
Leveraging these properties we claim that for any ${\boldsymbol{\mu}}\in\mathcal{D}$, $\lim_{n\to\infty}\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})=\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$. Since we have already proven uniqueness of $\boldsymbol{\lambda}$} in statement 2, we say $\hat{{\boldsymbol{\lambda}}}_{n}$ uniquely minimises $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Now from the $(\epsilon,\delta)$-definition of limits we say if $\Lambda_{\hat{\mathcal{F}}_{n}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is an $\epsilon$-cover of $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$ for $\epsilon>0$, then $|\hat{{\boldsymbol{\lambda}}}_{n}-{\boldsymbol{\lambda}}^{*}|\leq\delta$ for $\delta>0$ sufficiently small. It implies for a sequence of $\{\hat{\lambda_{n}}\}_{n\in\mathbb{N}}$ we claim $\lim_{n\to\infty}\hat{{\boldsymbol{\lambda}}}_{n}={\boldsymbol{\lambda}}^{*}$ i.e the sequence convergence. Therefore we conclude by the statement itself
@@ Appendix E Characterization of the Unique Optimal Policy: Proof of Theorem 3
Statement 1: Convexity of feasible space and optimal set function. Let us first analyse the properties of $\hat{\mathcal{F}}$. For any two member of $\boldsymbol{\omega}_{1},\boldsymbol{\omega}_{2}\in\hat{\mathcal{F}}$ satisfying $\tilde{A}\boldsymbol{\omega}_{1}\leq 0$ and $\tilde{A}\boldsymbol{\omega}_{2}\leq 0$, their convex combination for any $\alpha\in[0,1]$,
$\tilde{A}\left(\alpha\boldsymbol{\omega}_{1}+(1-\alpha)\boldsymbol{\omega}_{2}\right)=\alpha\tilde{A}\boldsymbol{\omega}_{1}+(1-\alpha)\tilde{A}\boldsymbol{\omega}_{2}\leq 0$
@@
(a) The complement of $\hat{\mathcal{F}}$, $\hat{\mathcal{F}}^{c}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}:\tilde{A}\boldsymbol{\pi}>0\}$ is an open set. (b) we have already proven the limit of $\hat{\mathcal{F}}$ to be $\mathcal{F}$ which is always contained by $\hat{\mathcal{F}}$.
@@
[…]
@@
Statement 4: Upper hemi-continuity of optimal allocation function. We refer to (Magureanu et al.,, 2014) (see Lemma 12) for this proof. We denote $Q(\tilde{A}^{\prime})\mathrel{\triangleq}\lim_{\hat{\mathcal{F}}\to\mathcal{F}}\max_{\boldsymbol{\omega}\in\Delta_{K}}\bigg\{\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}^{\prime\prime}\boldsymbol{\omega}\bigg|\tilde{A}^{\prime\prime}\boldsymbol{\omega}\leq 0,\boldsymbol{\omega}_{a}\geq 0\forall i\in[{K}]\bigg\}=\boldsymbol{\omega}({\boldsymbol{\mu}})$ where $\tilde{A}^{\prime\prime}\in\mathbb{R}^{{K}\times{K}}$ is the rank-1 update of $\tilde{A}^{\prime}$ which is a sub-matrix of $\tilde{A}$ with K{K} number of active constraints. We define limiting set as
$Q^{\star}(\tilde{A}^{\prime\prime})=\bigg\{\boldsymbol{\omega}:\lim_{\hat{\mathcal{F}}\to\mathcal{F}}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)=Q(\tilde{A}^{\prime\prime})\bigg|\tilde{A}^{\prime\prime}\boldsymbol{\omega}\leq 0,\boldsymbol{\omega}_{a}\geq 0\forall i\in[{K}]\bigg\}=\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$
@@ Appendix F Lagrangian Lower Bound for Gaussians: Proof of Theorem 4
$T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})=\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{2\sigma^{2}\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}\,.$
@@
[…]
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\lambda}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$
@@
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\gamma\in\mathbb{R}_{+}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\sum_{a=1}^{K}\boldsymbol{\omega}_{a}\frac{({\boldsymbol{\mu}}_{a}-{\boldsymbol{\lambda}}_{a})^{2}}{2\sigma^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\lambda}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$ (9)
@@
$\mathcal{L}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\gamma\in\mathbb{R}_{+}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{-\sum_{a=1}^{K}\gamma^{2}\frac{\boldsymbol{v}_{a}^{2}\sigma^{2}}{2\boldsymbol{\omega}_{a}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}-\gamma\left(\sum_{a=1}^{K}{\boldsymbol{\mu}}_{a}\boldsymbol{v}_{a}-r\right)\right\}$ (10)
@@
[…]
$=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\left\{\frac{1}{2\sigma^{2}}\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\right\}$